# Towards Excitations and Dynamical Quantities in Correlated Lattices with Density Matrix Embedding Theory

### v1

@@ Acknowledgements.
This project is sponsored by the US Department of Energy, Office of Science, Basic Energy Sciences, through Award No. DE-SC0018140. We thank Zhihao Cui and Shunyue Yuan for the development of the libdmet2 package that the excited-state code is based on. We thank Runze Chi and Yuhang Ai for useful discussions.
@@ Appendix A Practical Estimation of Hamiltonian matrices by Democratic Partitioning
We evaluate the transition density matrices with Eqs. 10 and 11 to illustrate the principle of democratic partitioning. Consider the following Hamiltonian:
@@
Eq. 24 can be thought of as obtained from an effective second-quantized Hamiltonian $\tilde{H}_{x}$ in the embedding space $A_{x}+B_{x}$:
@@ Appendix B Specification of the generalized excitation bases
In section III.0.2, we use a generalized excitation basis (denoted as PG in that context) truncated from the full general excitation basis. For neutral excitations, the excitation basis contains single-site projectors $\hat{P}_{s_{i}}$ as well as the spin-flip operators $\hat{s}_{\pm}$. Destruction operators $\hat{a}_{\alpha},\hat{a}_{\beta}$ are further included in addition to $\hat{P}_{s_{i}}$, $\hat{s}_{\pm}$, when computing the single-particle Green’s function.

### v2

@@ Acknowledgements.
This project is sponsored by the US Department of Energy, Office of Science, Basic Energy Sciences, through Award No. DE-SC0018140. We thank Zhihao Cui and Shunyue Yuan for the development of the libdmet2 package that the excited-state code [46] is inspired by. We thank Runze Chi and Yuhang Ai for useful discussions.
Author Contribution Statements
  • Work conceptualization: G.K.C.(lead), C.L. (lead), S.L. (supporting)
  • Code implementation: C.L.(excitation energies), S.L.(generalized excitation operators, Green’s function and code refactoring), H.Z. (support for DMRG code)
  • Validation: S.L.(equal), C.L.(equal)
  • Data generation: S.L.(lead), C.L.(supporting)
  • Writing - original draft: S.L.(lead), C.L.(supporting)
  • Writing - review & editing: S.L.(lead), G.K.C.(lead), C.L.(equal), H.Z.(equal)
@@ Appendix A Practical Estimation of Hamiltonian matrices by Democratic Partitioning
We evaluate the transition density matrices with Eqs. 9 and 10 to illustrate the principle of democratic partitioning. In practice, we firstly use the partitioning rules on the lattice Hamiltonian, then the resulting operators are contracted by the excitation bases to generate the effective Hamiltonian and overlap matrix in Eq. 6.
Consider the following Hamiltonian:
@@
Eq. 24 can be thought of as obtained from an effective second-quantized Hamiltonian $\tilde{H}_{x}$ ($E_{x}=\langle\Psi_{I}^{\mathrm{emb},x}|\tilde{H}_{x}|\Psi_{J}^{\mathrm{emb,x}}\rangle$) in the embedding space $A_{x}+B_{x}$:
@@ Appendix B Specification of the generalized excitation bases
In section III.2, we use a generalized excitation basis (denoted as PG in that context) truncated from the full general excitation basis. For neutral excitations, the excitation basis contains single-site projectors $\hat{P}_{s_{i}}$ as well as the spin-flip operators $\hat{s}_{\pm}$. Destruction operators $\hat{a}_{\alpha},\hat{a}_{\beta}$ are further included in addition to $\hat{P}_{s_{i}}$, $\hat{s}_{\pm}$, when computing the single-particle Green’s function.
@@
Supplementary Information
S1 Quantitative analysis of threshold
We test a broader range of overlap matrix eigenvalue cutoffs and observe its influence on the charged excited states and spectral function $A(k,\omega)$ in Fig. S2 and Fig. S3, and count the number of eigenvalues beyond the upper bound of $A(k,\omega)$ as a rough measurement of the quantity of unphysical eigenstates, shown in Fig. S1.
Both the $U=8$ and $U=4$ lines of Fig. S1 show the same trend of a decrease in the number of unphysical eigenstates: The number of unphysical states drops rapidly before some threshold (0.1 for $U=8$ and 0.04 for $U=4$), then it slowly decreases after further increasing the threshold. The profile of the spectra slightly change before threshold $0.04$ for $U=4$, and then the spectral shape is stable, while for $U=8$ shown in Fig. S2, the shape of the spectra develops artifacts at larger thresholds. Nonetheless, the choice of overlap cutoff threshold remains empirical in practice.
Figure S1: The number of effective Hamiltonian eigenvalues beyond the upper bound of $A(k,\omega)$ for different overlap eigenvalue cutoffs.
Figure S2: $A(k,\omega)$ of the Hubbard model $L=30,U=8$ with a larger range of overlap cutoff threshold ($0.01\sim 0.50$).
Figure S3: $A(k,\omega)$ of the Hubbard model $L=30,U=4$ with a larger range of overlap cutoff threshold ($0.01\sim 0.50$).
S2 Dependence on patch sizes of Green’s functions
We illustrate the dependence on patch size with additional computation using two-site patches in Figs. S4 and S5 using the excitation operator ansatz $P$ and $PG$ listed in Table S1 constructed in the same way as for three-site patches (Table 2 and 3. We clearly observe the improvement of the $S(k,\omega)$ structural factor from the two-site $P/PG1$ ansatz to the three-site $P/PG$ ansatz, especially at $\dfrac{\pi}{2}<k<\pi$ where the dispersion of the three-site patch is closer to that of the Bethe Ansatz and tDMRG. However the improvement is harder to see in the charge structure factor $N(k,\omega)$.
Type Operators Count
P $\hat{P}\otimes\hat{P}$ 16
PG1 $\hat{P}\otimes\hat{P},\hat{s}_{+}\otimes\hat{s}_{-},\hat{s}_{-}\otimes\hat{s}_{+}$ 18
Table S1: 2-site operator sites for calculating neutral excitations. $\hat{P}\in\{\hat{n}_{\alpha}\hat{n}_{\beta},\hat{n}_{\alpha}(1-\hat{n}_{\beta}),(1-\hat{n}_{\alpha})\hat{n}_{\beta},(1-\hat{n}_{\alpha})(1-\hat{n}_{\beta})\}$.
Figure S4: Spin dynamical structure factor $S(k,\omega)$ of the 30-site 1D Hubbard model from DMET (two-site and three-site patches), benchmark tDMRG results and Bethe Ansatz.
Figure S5: Charge dynamical structure factor $N(k,\omega)$ of the 30-site 1D Hubbard model from DMET (two-site and three-site patches), benchmark tDMRG results and Bethe Ansatz.